# Generalized Pulse Design in Floquet Engineering: Application to Interacting Spin Systems

### v1

@@ Appendix D Example Physical Systems
We briefly review how Alkali-based polar molecules, e.g., CsRb, and Rydberg atoms lead to effective spin models. Such experiments are captured by [47, 33, 34, 31, 35, 36, 32] Eqs. (1), (4), and (5) because in both cases, particles can be trapped in optical tweezers or optical lattices [32, 8, 48, 10, 12, 49, 50]. Localization of the particles into well-defined sites and mapping to pseudospin degrees of freedom implies that these systems are captured by spin lattice models.
Polar molecules in certain rotational states interact due to their dipole moment, $d_{0}$. Modeling the molecules as dipole moments in a planar array, e.g., a square lattice, we can define a quantization axis along a symmetry-breaking constant (DC) electric field pointing perpendicular to the plane. To define a pseudospin one can restrict to the subspace of rotational states where $\mathcal{N}\in\{0,1\}$ is the total rotational angular momentum and $m_{\mathcal{N}}=0$ is the $z$-component of the angular momentum [47].
[…]
@@
$a$ is the distance between nearest neighbors in the two-dimensional lattice, and the site-dependence of the interaction has a dipolar form, $\mathcal{V}_{j,j^{\prime}}=|j-j^{\prime}|^{-3}$. The relative interaction strengths can be characterized by an angle $\theta=\tan^{-1}(J_{\perp}/J_{z})$ that is tunable with the external electric field strength. The domain of $\theta$ for these physical systems ranges from $\pi/2$ to about $\pi/9$ so that the relative strengths cannot be negative [35]. This shows that a simple definition of a pseudospin using just rotational states of Alkali-based polar molecules leads to a sign restriction in $J_{\perp}$ and $J_{z}$. The restriction in turn limits the ability for global pulse protocols to convert a dipolar XXZ model into a purely Ising model. It would be interesting to explore the possibility of interleaving pulse schemes [44] designed to eliminate the long-range tail of the dipolar interaction to engineer bipartite models for molecules.
Recent experiments with Rydberg atoms have been able to engineer XY models and study their tunability [9, 10, 11]. Work studying circular Rydberg atoms trapped in optical tweezer arrays [43] also argues for the ability to realize dipolar XXZ models where a relative minus sign between the XY and ZZ terms can be obtained [43]. Consequently circular Rydberg atom systems can be converted from a dipolar XXZ model to a purely Ising model using global pulsing.

### v2

@@ Appendix D Example Physical Systems
We briefly review how Alkali-based polar molecules, e.g., CsRb, and Rydberg atoms lead to effective spin models. Such experiments are captured by [DeMille2002, Barnett2006, WALL2010a, Saffman2010, muller2010, Wall2014, NGUYEN2018] Eqs. (1), (4), and (5) because in both cases, particles can be trapped in optical tweezers or optical lattices [NGUYEN2018, BLUVSTEIN2021a, EBADI2021b, scholl2022microwave, CHRISTAKIS2023b, Holland2023, Bao2023]. Localization of the particles into well-defined sites and mapping to pseudospin degrees of freedom implies that these systems are captured by spin lattice models.
Polar molecules in certain rotational states interact due to their dipole moment, $d_{0}$. Modeling the molecules as dipole moments in a planar array, e.g., a square lattice, we can define a quantization axis along a symmetry-breaking constant (DC) electric field pointing perpendicular to the plane. To define a pseudospin one can restrict to the subspace of rotational states where $\mathcal{N}\in\{0,1\}$ is the total rotational angular momentum and $m_{\mathcal{N}}=0$ is the $z$-component of the angular momentum [DeMille2002].
A dipolar XXZ model arises due to Stark shift in the energy levels caused by the electric field. In that case, $|\uparrow\rangle$ is the Stark-shifted $|\mathcal{N}=0,m_{\mathcal{N}}=0\rangle$ state and $|\downarrow\rangle$ the Stark-shifted $|\mathcal{N}=1,m_{\mathcal{N}}=0\rangle$ state. In so doing, we obtain values for $J_{\perp}$ and $J_{z}$ that depend on the transition amplitudes of the Stark-shifted states and $d_{0}$. One can then show that Eq. (4) offers an excellent approximation where [Barnett2006, muller2010]:
@@
$a$ is the distance between nearest neighbors in the two-dimensional lattice, and the site-dependence of the interaction has a dipolar form, $\mathcal{V}_{j,j^{\prime}}=|j-j^{\prime}|^{-3}$. The relative interaction strengths can be characterized by an angle $\theta=\tan^{-1}(J_{\perp}/J_{z})$ that is tunable with the external electric field strength. The domain of $\theta$ for these physical systems ranges from $\pi/2$ to about $\pi/9$ so that the relative strengths cannot be negative [muller2010]. This shows that a simple definition of a pseudospin using just rotational states of Alkali-based polar molecules leads to a sign restriction in $J_{\perp}$ and $J_{z}$. The restriction in turn limits the ability for global pulse protocols to convert a dipolar XXZ model into a purely Ising model. It would be interesting to explore the possibility of interleaving pulse schemes [Lee2016] designed to eliminate the long-range tail of the dipolar interaction to engineer bipartite models for molecules.
Recent experiments with Rydberg atoms have been able to engineer XY models and study their tunability [GEIER2021c, scholl2022microwave, ZHAO2023]. Work studying circular Rydberg atoms trapped in optical tweezer arrays [NGUYEN2018] also argues for the ability to realize dipolar XXZ models where a relative minus sign between the XY and ZZ terms can be obtained [NGUYEN2018]. Consequently, circular Rydberg atom systems can be converted from a dipolar XXZ model to a purely Ising model using global pulsing.